Solve the equation 7cot⁡θ=tan⁡θ for all nonnegative values of θ less than 360∘ . Do by calculator, if needed, and give the answers to three significant digits in the order of increasing.
Answer by stanbon(75887)

Solve the equation 7cot⁡θ=tan⁡θ for all nonnegative values of θ less than 360∘ . Do by calculator, if needed, and give the answers to three significant digits in the order of increasing.
--------------------
tan(t)/cot(t) = 7
tan^2(t) = 7
----
tan(t) = sqrt(7)
t = 69.300 or t = 180+69.300 = 249.300 degrees
------------
Cheers,
Stan H.
----------